OpenStudy (anonymous):

Write the equation of the line in slope-intercept form with a slope of -6 and a y-intercept of 9.

jimthompson5910 (jim_thompson5910):

the basic slope intercept form equation is y = mx+b

jimthompson5910 (jim_thompson5910):

m is the slope b is the y-intercept

jimthompson5910 (jim_thompson5910):

the given slope is -6, so m = -6

jimthompson5910 (jim_thompson5910):

the y-intercept is 9, which is really the point (0,9) so b = 9

jimthompson5910 (jim_thompson5910):

I'm sure you see what to do from here

OpenStudy (anonymous):

is the answer y=-6x+9

jimthompson5910 (jim_thompson5910):

exactly

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

thank you!

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

can you help me with another?

jimthompson5910 (jim_thompson5910):

sure

jimthompson5910 (jim_thompson5910):

what's your question

OpenStudy (anonymous):

Choose the equation of the line passing through the point (6, -2) and parallel to y = x + 7. y = x + 8 y = x - 4 y = x + 4 y = x - 8

jimthompson5910 (jim_thompson5910):

parallel lines have equal slopes

jimthompson5910 (jim_thompson5910):

the slope of y = x+7 is 1 so the slope of our answer will also be 1

jimthompson5910 (jim_thompson5910):

so m = 1

jimthompson5910 (jim_thompson5910):

y = mx + b y = 1x + b now plug in the point you want this parallel line to go through so plug in (6,-2) or x = 6 and y = -2 to get -2 = 1(6) + b -2 = 6 + b solve for b

OpenStudy (anonymous):

wait... how do i do that?

jimthompson5910 (jim_thompson5910):

Subtract 6 from both sides.

jimthompson5910 (jim_thompson5910):

to isolate b

OpenStudy (anonymous):

i got -8

jimthompson5910 (jim_thompson5910):

same here

jimthompson5910 (jim_thompson5910):

m = 1 b = -8

jimthompson5910 (jim_thompson5910):

y = mx + b y = 1x - 8 y = x - 8

OpenStudy (anonymous):

thank you!

jimthompson5910 (jim_thompson5910):

sure thing